A first order reaction has a rate constant of 0.925 at 25 oC. Given that the activation energy is 34.6 kJ/mol, calculate the rate constant at 35 oC. (enter answer to 3 decimal places)
Answer #1

ln ( k2/k1) = ( E/R)(1/T1 -1/T2)

ln ( k2/0.925) = ( 34.6 x 1000 /8.314) ( 1/ 298-1/308)

k2 = 1.456